# Mixture of Experts Using Tensor Products

### v1

@@ 5.1 Backbone, Datasets and Evaluation
To evaluate the generalization capabilities of our models, we adopt the same benchmarking strategy as (tfew), utilizing a subset of tasks designated as held-out from the multitask training. This benchmark encompasses a diverse array of tasks, including sentence completion (COPA (roemmele2011choice), H-SWAG(zellers2019hellaswag) and Story Cloze (sharma2018tackling) datasets), natural language inference (ANLI (nie2019adversarial), CB (de2019commitmentbank) and RTE (dagan2005pascal)), coreference resolution (WSC (levesque2012winograd), Winogrande(sakaguchi2021winogrande)), and word sense disambiguation (WIC (pilehvar2018wic)). For each task, our evaluation strategy involves constructing sets of five few-shot training examples, which are generated by sampling subsets from each dataset using different seeds. We then report the median performance. It is note that the prompt examples from each dataset using the prompt templates from P3 (bach2022promptsource), using a randomly-sampled prompt template for each example.
@@
For the evaluation of our models, we employ the rank classification methodology as outlined by the tfew study. This approach involves ranking the model’s log-probabilities for all possible label strings associated with each task. The model’s prediction is deemed correct if the label string with the highest log-probability ranking corresponds to the correct answer. This method allows for a nuanced assessment of the model’s predictive accuracy by examining its ability to prioritize the correct label over others based on their calculated log-probabilities, offering a precise measure of its understanding and processing of the task at hand.
@@ 5.2 Baselines
In our comparative analysis, we initially set the benchmark by evaluating the performance of the TLoRA model against the traditional full fine-tuning approach, referred to as FullFT. In the FullFT scenario, we do not freeze any parameters of the pre-trained model, nor do we insert any adapters, allowing for a comprehensive update of the model’s parameters during fine-tuning. Subsequently, we contrast our method against a suite of established parameter-efficient fine-tuning (PEFT) baselines to study the efficiency and effectiveness of each in terms of training parameter utilization. These baselines include:
  • Adapter, as introduced by houlsby2019parameter, which involves inserting trainable layers while keeping the pre-trained model’s parameters fixed.
  • BitFit by zaken2021bitfit, which only fine-tunes the bias terms within the model.
  • LoRA proposed by lora, adjusting the low-rank adaptations of the weight matrices.
  • Compacter and Compacter++ by karimi2021compacter, which extend the adapter methodology with compact and efficient training strategies.
  • Prompt tuning (lester2021power) and Prefix tuning li2021prefix add some tunable vectors to the input or hidden layer of the base model.
  • FishMask by sung2021training, identifying and training a subset of parameters.
  • Intrinsic SAID as described by aghajanyan2020intrinsic, focusing on intrinsic sparse activations.
  • IA3 (tfew), emphasizing adaptability and efficiency.
We undertake a comparative analysis between routing approaches TensorPoly, Poly and employing a single expert mechanism without routing, specifically TLoRA and LoRA. This comparison aims to evaluate the impact of routing techniques on model performance and efficiency. By contrasting these models, we seek to understand how the dynamic allocation of tasks to specific experts in TensorPoly and Poly compares to the single expert models without shared information across tasks.
@@ 5.3 Results
Method Pre-Training Fine-Tuning
FullFT $d\times d$ $d\times d$
LoRA $2\times d\times r$ $2\times d\times r$
TLoRA $2\times N\times r\times\lceil\sqrt[N]{d}\rceil\times R$ $2\times N\times r\times\lceil\sqrt[N]{d}\rceil\times R$
Poly $2\times d\times r\times|S|+|\mathcal{T}|\times|S|$ $2\times d\times r\times|S|+|S|$
TensorPoly-I $2\times N\times r\times\lceil\sqrt[N]{d}\rceil\times R+|\mathcal{T}|\times R$ $2\times N\times r\times\lceil\sqrt[N]{d}\rceil\times R$ + $R$
TensorPoly-II $2\times N\times r\times\lceil\sqrt[N]{d}\rceil\times R+\mathcal{T}\times R\times N$ $2\times N\times r\times\lceil\sqrt[N]{d}\rceil\times R$ + $R\times N$
Table 2: Number of parameters (per layer) used for each method. $d$ is the input and output dimension of the training parameters. We assume they are identical. $r$ is the rank in the LoRA, where $r\ll d$. $N$ and $R$ are the order and rank of entangled tensor respectively. $S$ is the number of modules in Poly.
Table. 1 presents the mean downstream accuracy for 11 held-out tasks in the T0 benchmark. Notably, we did not train all these PEFT approaches, all the results are reported from (tfew). When evaluating the performance of various PEFT approaches against single expert performances, it is observed that many PEFT strategies achieve similar outcomes while utilizing a significantly smaller subset of training parameters compared to the FULLFT (Full Fine-Tuning) method. Additionally, the implementation of Low-Rank Adaptation (LoRA) within our training framework further illustrates the efficiency of these methods. Specifically, TLoRA achieves a competitive score of 66.1, closely trailing the original LoRA’s score of 66.8, while requiring only about 60% of the training parameters used by LoRA. This demonstrates that TLoRA not only matches the effectiveness of LoRA in terms of performance but also surpasses it in terms of parameter efficiency.
In our analysis, we initially contrast the modular model against the dense model, followed by a comparison of routing-based approaches with a single adapter strategy. Within this context, we utilize both LoRA and TLoRA as baselines for these routing techniques. According to the results presented in Table 1, the Poly model demonstrates superior performance over LoRA by a margin of 2.0 points. Moreover, TensorPoly exhibits an improvement of 3.2 points over the base TLoRA model, underscoring the efficacy of routing in enhancing multi-task generalization within the realm of multi-task transfer learning.
When evaluating various Poly variants, TensorPoly-I stands out by not only surpassing recent state-of-the-art achievements but also by outperforming TensorPoly-II, despite the latter employing a more granular routing function. This finding is particularly noteworthy, as it suggests that the increased specificity of the routing function in TensorPoly-II does not necessarily translate to superior performance. We will discuss this in Section 5.5.
@@ 5.3.1 Rank and Order Analysis
Figure 3: Explore the tensor Rank in the Entangled Tensor. We set order $N=2$ in this setting.
The rank of an entangled tensor serves as a measure of its capacity, denoting the fewest number of rank-one tensors required for its original tensor (§3.2). We focus on examining the relationship between the rank of tensor and the associated training parameters. Table 2 illustrates how the tensor rank $R$ and tensor order $N$ impact the selection of training and fine-tuning parameters. Illustrated in Figure 3, when the tensor rank is set to 1, it corresponds to the original TLoRA configuration. An increase in tensor rank necessitates a larger number of training parameters. Concomitantly, there is a notable enhancement in performance. This progression underscores the direct correlation between the tensor rank and the model’s capability.
Then tensor order $N$, correlates with the granularity of training parameters; a tensor of order 4 yields a finer-grained module compared to one of order 2. Our research examines the impact of varying tensor orders on performance outcomes. For simplicity, we constrain our analysis to tensors of order 2 and 4. Our results demonstrate that a tensor of order 2 outperforms one of order 4 by a margin of 1.7 for the TensorPoly-I. For TensorPoly-II, an order-2 tensor exceeds the performance of an order-4 tensor by 4.8, suggesting a balance must be struck between the size of experts and its efficacy. While higher-order tensors may conserve training parameters, this comes at the cost of diminished performance.
Model Adaptation Params ACC
LoRA[r=1] 2.2M 66.8
TLoRA [r=16] 1.4M 66.1
Poly[r=1] 17 M 68.8
TensorPoly-I[N=2] 12.2M 68.7
TensorPoly-I[N=4] 4.3M 66.9
TensorPoly-II[N=2] 13.3M 67.9
TensorPoly-II[N=4] 7.6M 63.1
Table 3: Results on T0 dataset, with T0-3B. The Adaptation parameter is the number of parameters in the finetuning process.
Model Adaptation Params ACC
LoRA[r=1] 2.2M 66.8
TLoRA [r=16] 1.4M 66.1
Poly-$\mu$[r=1] 2.2M 68.8
@@
TensorPoly-I-$\mu$[N=2] 1.4M 69.0
TensorPoly-II-$\mu$[N=2] 1.4M 68.2
Table 4: Finetuning the modules only. The Adaptation parameter is the number of parameters in the finetuning process.
5.4 Routing Analysis
@@
Figure 4: Distribution of experts weights in the T0 model. X-axis is the 32 experts in each layer (For Q,K,V,O, each corresponds to 8 experts). Y-axis denotes the 24 transformer layers in T0 model. We show routing distributions for different kinds of tasks.
In the paper referenced as mhr, the MHR study demonstrates that fine-tuning solely the routing function can yield competitive outcomes. This insight provides a valuable perspective for our investigation into various routing strategies within the TensorPoly framework. In line with this approach, we focus on fine-tuning exclusively the routing function during the few-shot adaptation process, indicated by the notation -$z$. This methodological choice allows us to isolate the impact of the routing function’s optimization on the overall performance of the TensorPoly model, thereby offering a clearer understanding of how dynamic routing contributes to the adaptability and efficiency of the model in few-shot learning scenarios.
As detailed in Table 5, an initial observation reveals that the routing parameters necessitate a remarkably small number of training parameters. The performance metrics of Poly-$z$, TensorPoly-I-$z$, and TensorPoly-II-$z$ lag behind their counterparts where both the modules and the routing function undergo fine-tuning. This disparity highlights the critical role that modules play in the fine-tuning process for TensorPoly. The findings suggest that while optimizing the routing function alone can contribute extreme parameter efficiency, the integration of module fine-tuning is indispensable for achieving the best possible performance. This underlines the synergy between modules and routing functions in the TensorPoly architecture.
Model Adaptation Params ACC $\Delta$
Poly-$z$ 3.5k 65.4 -3.4
MHR-$z$ 220K 68.3 -0.9
TensorPoly-I-$z$[N=2] 3.5k 64.3 -4.4
TensorPoly-I-$z$[N=4] 3.5k 62.7 -0.4
TensorPoly-II-$z$[N=2] 6.9k 62.5 -5.4
TensorPoly-II-$z$[N=4] 13.8k 60.3 -2.8
Table 5: Results on T0 dataset, with T0-3B. We only fine-tuning the routing function.
5.5 Discussion
In modular language models, tensor product routing is poised to construct more intricate patterns. This capability arises from its method of aggregating activated modules through tensor products, as opposed to the conventional linear routing approach. However, our experimental results diverged from theoretical expectations, revealing that tensor product routing did not contribute to improved final performance in our models. This discrepancy prompts a future line of inquiry: we plan to explore whether there exist specific benchmarks or conditions under which tensor product routing could demonstrate its purported benefits. Identifying such scenarios will be crucial for harnessing the potential advantages of tensor product routing in modular language models.
6 Conclusion
We introduce a novel modular language model named TensorPoly. This model incorporates tensorized modules, specifically TLoRA, to significantly reduce the number of training parameters required by the traditional LoRA approach. We employ two distinct strategies for aggregating activated modules: a linear routing function, referred to as TensorPoly-I, and a more finely trained tensor product routing, named TensorPoly-II. Our evaluation across various multi-task learning scenarios reveals that modular language models, such as TensorPoly, surpass the performance of single-adapter models. This underscores the importance of sharing task information through a routing function in multi-task learning contexts. Notably, TensorPoly-I achieves state-of-the-art results, highlighting the effectiveness of the TensorPoly framework. However, TensorPoly-II does not outperform TensorPoly-I in our experimental settings, suggesting areas for further investigation in future research.
Broader Impact Statement
Author Contributions
If you’d like to, you may include a section for author contributions as is done in many journals. This is optional and at the discretion of the authors. Only add this information once your submission is accepted and deanonymized.
Acknowledgments
Anonymized for double-blind review.
Appendix A Appendix
A.1 FLOP Analysis
Consider the tensor product of a vector $a$ of size $n\times 1$ with a vector $b$ of size $1\times n$, resulting in a matrix $C$ of size $n\times n$. The product operation involves $n^{2}$ multiplications due to the $n$ elements in each of the vectors $a$ and $b$. As described in Table 2 and shown in Figure 2, the additional computational effort in TLoRA stems from the tensor product operations, dictated by the order of the tensor. Specifically, for parameterizing with a dimension $d$, rank $r$, and tensor rank $R$, the number of extra computations required is approximately $d\times r\times R$. There may add some time consumption analysis
N=1,R=1
N=1,R=2
N=1,R=3
N=1,R=4
N=1,R=5
N=1,R=6
N=1,R=7
N=1,R=8
N=2,R=1
N=1,R=2
N=1,R=3
N=1,R=4
N=1,R=5
N=1,R=6
N=1,R=7
N=1,R=8
Figure 5: Weights distributions across different experts. We set N=2, R=8.

### v2

@@ 5.1 Datasets and Evaluation
To evaluate the generalization capabilities of our models, we adopt the same benchmarking strategy as (Liu et al., 2022), utilizing a subset of tasks designated as held-out from the multitask training. This benchmark encompasses a diverse array of tasks, including sentence completion (COPA (Roemmele et al., 2011), H-SWAG(Zellers et al., 2019) and Story Cloze (Sharma et al., 2018) datasets), natural language inference (ANLI (Nie et al., 2019), CB (De Marneffe et al., 2019) and RTE (Dagan et al., 2005)), coreference resolution (WSC (Levesque et al., 2012), Winogrande (Sakaguchi et al., 2021)), and word sense disambiguation (WIC (Pilehvar & Camacho-Collados, 2018)). For each task, our evaluation strategy involves constructing sets of five few-shot training examples, which are generated by sampling subsets from each dataset using different seeds. We then report the median performance. It should be noted that the prompt examples from each dataset using the prompt templates from P3 (Bach et al., 2022).
@@
For the evaluation of our models, we employ the rank classification methodology as outlined by the Liu et al. (2022) study. This approach involves ranking the model’s log-probabilities for all possible label strings associated with each task. The model’s prediction is deemed correct if the label string with the highest log-probability ranking corresponds to the correct answer. This method allows for a nuanced assessment of the model’s predictive accuracy by examining its ability to prioritize the correct label over others based on their calculated log-probabilities, offering a precise measure of its understanding and processing of the task at hand.
@@ 5.2 Baselines
Method Pre-Training Fine-Tuning
FullFT $d^{2}$ $d^{2}$
LoRA $2dr$ $2dr$
TLoRA $2Nr\lceil\sqrt[N]{d}\rceil R$ $2Nr\lceil\sqrt[N]{d}\rceil R$
Poly $2dr|S|+|\mathcal{T}||S|$ $2dr|S|+|S|$
TensorPoly-I $2Nr\lceil\sqrt[N]{d}\rceil R+|\mathcal{T}|R$ $2Nr\lceil\sqrt[N]{d}\rceil R+R$
TensorPoly-II $2Nr\lceil\sqrt[N]{d}\rceil R+|\mathcal{T}|RN$ $2Nr\lceil\sqrt[N]{d}\rceil R+RN$
Table 2: Number of parameters (per layer) used for each method. $d$ is the input and output dimension of the training parameters. We assume they are identical. $r$ is the rank in the LoRA, where $r\ll d$. $N$ and $R$ are the order and rank of entangled tensors respectively. $S$ is the number of modules in Poly.
In our comparative analysis, we initially set the benchmark by evaluating the performance of the TLoRA model against the traditional full fine-tuning approach, referred to as FullFT. To facilitate a fair comparison with baseline methodologies, we have chosen the T0-3B model, consistent with the approach described in the IA3 paper by Liu et al. (2022). In the FullFT scenario, we do not freeze any parameters of the pre-trained model, nor do we insert any adapters, allowing for a comprehensive update of the model’s parameters during fine-tuning. Subsequently, we contrast our method against a suite of established parameter-efficient fine-tuning (PEFT) baselines to study the efficiency and effectiveness of each in terms of training parameter utilization. These baselines include: Adapter, as introduced by Houlsby et al. (2019b), which involves inserting trainable layers while keeping the pre-trained model’s parameters fixed; BitFit by Zaken et al. (2021), which only fine-tunes the bias terms within the model; LoRA proposed by Hu et al. (2021), adjusting the low-rank adaptations of the weight matrices; Compacter and Compacter++ by Karimi Mahabadi et al. (2021), which extend the adapter methodology with compact and efficient training strategies; Prompt tuning (Lester et al., 2021) and Prefix tuning Li & Liang (2021) add some tunable vectors to the input or hidden layer of the base model; FishMask by Sung et al. (2021), identifying and training a subset of parameters; Intrinsic SAID as described by Aghajanyan et al. (2020), focusing on intrinsic sparse activations; and IA3 (Liu et al., 2022), emphasizing adaptability and efficiency.
We perform a comparative analysis between routing approaches TensorPoly, Poly/MHR and dense models without routing, specifically TLoRA and LoRA. This comparison aims to evaluate the impact of routing techniques on model performance and efficiency. By contrasting these models, we seek to evaluate how the dynamic allocation of tasks to specific experts in TensorPoly and Poly/MHR compares to the dense models.
@@ 5.3 Results
Tab. 1 presents the mean downstream accuracy for 11 held-out tasks in the T0 benchmark. We reported most of the results from (Liu et al., 2022). When evaluating the performance of various PEFT approaches against single-expert performances, it is observed that many PEFT strategies achieve similar outcomes while utilizing a significantly smaller subset of training parameters compared to the FullFT method. Furthermore, the LoRA method within our training framework further illustrates the efficiency of these methods. Specifically, TLoRA achieves a competitive score of 66.1, closely trailing the original LoRA’s score of 66.8, while requiring only fewer training parameters used by LoRA. This demonstrates that TLoRA not only matches the effectiveness of LoRA in terms of performance but also achieves a higher parameter efficiency.
In our analysis, we initially contrast the modular model against the dense model, followed by a comparison of routing-based approaches with a single adapter strategy. Within this context, we utilize both LoRA and TLoRA as baselines for these routing techniques. According to the results presented in Tab. 1, the Poly model demonstrates superior performance over LoRA by a margin of 2.0 points. Moreover, TensorPoly exhibits an improvement of 3.2 points over the base TLoRA model, underscoring the superiority of routing in enhancing multi-task generalization within multi-task transfer learning.
When evaluating the various modular models, TensorPoly-I stands out by not only on par with the recent state-of-the-art achievements but also by outperforming TensorPoly-II, despite the latter employing a more granular routing function. This finding is particularly noteworthy, as it suggests that the increased specificity of the routing function in TensorPoly-II does not necessarily translate to superior performance. We will discuss this in Section 5.4.
@@ 5.3.1 Rank and Order Analysis
Figure 4: Rank analysis in the TensorPoly-I, Left denotes the average accuracy over 11 held-out tasks according to different rank. Right is the validation loss in the multi-task pre-training process.
For TensorPoly, rank $R$ and order $N$ correlate with the number of experts. Tab. 2 illustrates how the tensor rank $R$ and the tensor order $N$ affect the training and adaptation parameters. As illustrated in Figure 4 (left), when the tensor rank is set to 1, it corresponds to the original TLoRA configuration. An increase in tensor rank necessitates a larger number of training parameters. Consequently, there is a notable enhancement in performance. In addition, we investigate if the rank will give more capacity to the model. As depicted in Figure 4 (right), as the increase of the rank, we can get a lower validation loss score during the multi-task pre-training. This progression underscores the direct correlation between the tensor rank and the capability of the model.
The tensor order $N$, correlates with the granularity of training parameters (experts) (§$\S$4.2); a tensor of order 4 yields a finer-grained module compared to one of order 2. Our research examines the impact of varying the tensor orders on performance outcomes. For simplicity, we constrain our analysis to tensors of order 2 and 4. Our results demonstrate that a tensor of order 2 outperforms one of order 4 by a margin of 1.7 for the TensorPoly-I. For TensorPoly-II, an order-2 tensor exceeds the performance of an order-4 tensor by 4.8, suggesting that a balance must be struck between the granularity of experts and its efficacy. Although higher-order tensors may conserve training parameters, this comes at the cost of diminished performance.
Model R N Multi-Task Params Adaptation Params ACC
LoRA - - 2.2M 2.2M 66.8
TLoRA 1 2 1.4M 1.4M 66.1
Poly - - 17M 2.2M 68.8
TensorPoly-I 8 2 12.2M 1.4M 68.7
TensorPoly-I 20 2 27.8M 1.4M 69.3
TensorPoly-I 8 4 4.3M 1.4M 66.9
TensorPoly-II 8 2 13.3M 1.4M 67.9
TensorPoly-II 20 2 33.2M 1.4M 68.0
TensorPoly-II 8 4 7.6M 1.4M 63.1
Table 3: Rank $R$ and Order $N$ analysis. The Adaptation parameter is the number of parameters required to learn a new downstream task. Multi-Task params is the number of additional parameters that must be conserved after multi-task pre-training to enable transfer to a downstream task.
5.3.2 Routing Analysis
Model modules routing Adaptation Params ACC
LoRA - 2.2M 66.8
TLoRA - 1.4M 66.1
Modular LMs
Poly-$\mu$ 2.2M 68.8
@@ 5.3.2 Routing Analysis
TensorPoly-I-$\mu$ 1.4M 69.0
TensorPoly-II-$\mu$ 1.4M 68.2
Poly-$z$ 3.5k 65.4
MHR-$z$ 220K 68.3
TensorPoly-I-$z$ 8.6k 66.3
TensorPoly-II-$z$ 17.3k 65.4
Poly 2.2M 68.8
MHR 2.2M 69.2
TensorPoly-I 1.4M 69.3
TensorPoly-II 1.4M 68.0
Table 4: We compare several fine-tuning approaches. $-\mu$ represents we only fine-tune the modules. $-z$ means we only fine-tune the routing functions. We set the order $N=2$ for this comparison.
Routing Only
In the paper Caccia et al. (2023), the MHR study demonstrates that fine-tuning solely the routing function can yield competitive outcomes. This insight provides a valuable perspective for our investigation into various routing strategies within the TensorPoly framework. In line with this approach, we focus exclusively on fine-tuning the routing function during the few-shot adaptation process, indicated by the notation $z$. This methodological choice allows us to isolate the impact of the optimization of the routing function on the overall performance of the TensorPoly model, thereby offering a clearer understanding of how dynamic routing contributes to the adaptability and efficiency of the model in few-shot learning scenarios.
As detailed in Tab. 4, an initial observation reveals that the routing parameters necessitate a small number of training parameters, achieving extreme parameter efficiency. TensorPoly-I-$z$ achieves a accuracy score of 66.3 while only using 8.6k adaptation parameters. TensorPoly-II-$z$ achieves 65.4 with 17.3k parameters. MHR-$z$ achieve the 68.3 with 220K adaptation parameters. The results indicate that TensorPoly-I-$z$ can achieve competitive results as LoRA with higher parameter efficiency. Notably, the average accuracy of TensorPoly-I-$z$, and TensorPoly-II-$z$ lag behind their counterparts where both the modules and the routing function undergo fine-tuning. This disparity highlights the critical role that modules play in the fine-tuning process for TensorPoly.
Modules Only
To verify whether the routing is important in few-shot fine-tuning, we discard the routing function and average the pre-trained modules in the fine-tuning process, indicated by the notation $-\mu$. The result in Tab. 4 shows that there is only a slight decrease for MHR-$\mu$ and TensorPoly-I-$\mu$ compared to their counterparts where both the modules and the routing function are fine-tuned. This result is consistent with the conclusion in Caccia et al. (2023).
5.4 Discussion
5.4.1 Investigate more about the latent expert approach.
Since we add more training parameters (we use more experts), it is necessary to investigate if the improvement is caused by adding more training parameters. To investigate this, In Caccia et al. (2023), they assign a binary module allocation to each data point in a minibatch, disregarding task information. At test time, the learned modules are averaged into one single module. This approach is named random-$\mu$, the random-$\mu$ performs similar to single LoRA, which proved that the "routing" function is important for latent-expert approach in multi-task transfer learning.
Figure 5: Validation loss compare across different models.
5.4.2 Why TensorPoly-II underperform the TensorPoly-I?
Finer-grained routing has been shown to be effective in multi-head routing function (Caccia et al., 2023). However, our experimental results underscore that the finer-grained tensor product routing did not contribute to improved final performance in our models. This discrepancy prompts a future line of inquiry: we plan to explore whether there exist specific benchmarks or conditions under which the tensor-product interaction could demonstrate its purported benefits. Identifying such scenarios will be crucial to harnessing the potential advantages of tensor product routing in modular language models.
6 Conclusion
We introduce a novel modular language model named TensorPoly. This model incorporates tensorized modules, specifically TLoRA, to significantly reduce the number of training parameters required by the traditional LoRA approach. We employ two distinct strategies for aggregating activated modules: TensorPoly-I directs to each rank and a more finely routing, named TensorPoly-II targets each order of the tensor. Our evaluation across various multi-task learning scenarios reveals that modular language models, such as TensorPoly, surpass the performance of single-adapter models. This underscores the importance of sharing task information through a routing function in multi-task learning contexts. Notably, TensorPoly-I achieves state-of-the-art results, highlighting the effectiveness of the TensorPoly framework. However, TensorPoly-II does not outperform TensorPoly-I in our experimental settings, suggesting areas for further investigation in future research.
Broader Impact Statement
Acknowledgments
We acknowledge the support of Alessandro Sordoni for optimizing the code and discussion.
Appendix A Appendix
A.1 Dataset
For each task in T0 benchmark(Sanh et al., 2021), there are a few examples to help with the task adaptation. The details are described in Tab 5.
Dataset Val Num Few Shot
COPA 100 32
WIC 638 32
RTE 277 32
CB 56 32
Winograd 1267 50
RTE1 1000 32
RTE2 1000 32
RTE3 1200 32
WSC 104 32
H-swag 10042 20
Table 5: Dataset statistics of the T0 benchmark
A.2 FLOP Analysis
Model Trainable parameters Training time ACC
LoRA 2.2M 1d 14h 53m 27s 66.8
TLoRA 1.4M 2d 3h 13m 45s 66.1
Poly 17M 1d 16h 16m 43s 68.8
MHR 17M 1d 17h 5m 20s 69.2
TensorPoly-I (rank=8) 12.2M 2d 3h 13m 8s 68.7
TensorPoly-II (rank=8) 13.3M 2d 3h 48m 19s 67.9
TensorPoly-I(rank=20) 27.8M 2d 6h 6m 3s 69.3
TensorPoly-II(rank=20) 33.2M 2d 2h 55m 59s 68.0
Table 6: Training time analysis across different models.
Consider the tensor product of a vector $a$ of size $n\times 1$ with a vector $b$ of size $1\times n$, resulting in a matrix $C$ of size $n\times n$. The product operation involves $n^{2}$ multiplications due to the $n$ elements in each of the vectors $a$ and $b$. As described in Table 2 and shown in Figure 3, the additional computational effort in TLoRA stems from the tensor product operations, dictated by the order of the tensor. Specifically, for parameterizing with a dimension $d$, rank $r$, and tensor rank $R$, the number of extra computations required is approximately $d\times r\times R$.
To test the time consumption according to the Flop analysis, we show the training time in Tab 6. The results indicate that although the tensor product can reduce the training parameters, we need more training time according to the original LoRA adapters.
A.3 Few-shot finetuning parameter size over different PEFT approaches
We investigate the training parameter size over PEFT approaches in the few-shot finetuning process. As presented in Tab. 7, TLoRA obtains competitive results while using only 60% training parameters than the original LoRA in this train setting. Compared with the latent experts approach Poly, TensorPoly-I with tensor rank $R=8$ and order $N=2$, get almost the same results with fewer finetuning parameters. TensorPoly-I(R=20,N=2) obtain the best results with 27.8 finetuning parameters. Notably, the finetuning parameters are not the adaption parameters we present in Tab. 3.
Model Finetuning parameter size ACC
FullFT 3B 63.3
BitFiT(with LN) 1.3M 54.6
LayerNorm 250K 54.2
Adapter 12.9M 61.6
Compacter 807K 61.0
Compacter++ 540K 61.5
Prompt (10) 41K 50.9
Prompt (100) 409K 47.7
Prefix Tuning 576K 57.8
FishMask(0.2%) 6M 61.1
FishMask(0.02%) 600K 59.0
SAID 500K 57.5
LoRA 9.1M 62.8
IA3 540K 64.6
w/pre-train
LoRA 2.2M 66.8
TLoRA 1.4M 66.1
Poly 17M 68.8
TensorPoly-I(R=8,N=2) 12.2M 68.7
TensorPoly-I(R=20,N=2) 27.8M 69.3
TensorPoly-II(R=8,N=2) 13.3M 67.9
TensorPoly-II(R=20,N=2) 33.2M 68.0
Table 7: We show the parameters of different approaches in the finetuning process. For the baselines, we report the results from the paper (Liu et al., 2022).
A.4 Routing across different tasks
@@
Figure 6: The X-axis is the 32 experts in each layer (For Q,K,V,O, each corresponds to 8 experts). Y-axis denotes the 24 transformer layers in the T0 model. We show routing distributions for different kinds of tasks.
To investigate the routing function across different tasks, we plot the routing distributions for tasks such as Question Answering and Summarization. As depicted in Figure 6, the routing function distribution varies across different layers for each task. Notably, Closed-Book QA exhibits greater sparsity compared to multi-choice QA. We will conduct further analysis of the routing function in various tasks to gain deeper insights.
A.5 Latent Experts Visualization
We visualize the latent experts in our models. As depicted in Figure 7, we draw the expert’s weight distribution in layer [0,4,16,23]. The distribution of expert weights varies in different layers. Especially, The expert weights in layer 16 seem completely different from the experts in layers [0, 4, 23]. We also draw the similarity of expert weights across different layers in Figure 8.
Layer 0, N=1,R=1
Layer 0, N=1,R=8
Layer 0, N=2,R=1
Layer 0, N=2,R=8
Layer 4, N=1,R=1
Layer 4, N=1,R=8
Layer 4, N=1,R=1
Layer 4, N=1,R=8
Layer 16, N=1,R=1
Layer 16, N=1,R=8
Layer 16, N=1,R=1
Layer 16, N=1,R=8
Layer 23, N=1,R=1
Layer 23, N=1,R=8
Layer 23, N=1,R=1
Layer 23, N=1,R=8
Figure 7: Weights distributions for latent experts across different layers. $N$ is the order and $R$ is the tensor rank. Each weight matrix here is $\mathcal{A}_{i,k}\in\mathbb{R}^{r\times\lceil\sqrt[N]{d}\rceil}$.
(a) Layer 0
(b) Layer 4
(c) Layer 16
(d) Layer 23
Figure 8: Expert similarity in each layer in TensorPoly-II. We compare the similarity of 16 experts in each layer. 0-7 refers to experts with order 1, 8-15 refers to the experts with order 2.